OpenStudy (cutiekawdfi):

help please will fan and medal (((: find the real roots of each equation by graphing. 1. x^2 + 3 = x^3 - 5

OpenStudy (faiqraees):

make a graph of y=x^2 + 3 then make a graph of y=x^3 - 5 The intersection points are the roots

OpenStudy (cutiekawdfi):

.-.

OpenStudy (faiqraees):

To make a graph randomly take x values like 1,2,3 and plug them in equation to get the value of y and you will have the x and y coordinate. Plot it accordingly and you will get a graph

OpenStudy (cutiekawdfi):

okay thank yu

OpenStudy (jv0408):

Find the solutions (roots) of a quadratic equation using the graphing calculator. Solve: Since this equation is set equal to zero, the roots will be the locations where the graph crosses the x-axis (if the roots are real numbers). (Remember that the x-axis is really just y = 0.)

OpenStudy (jv0408):

hope that helps